If you wish to increase the amount of current in a resistor from 100mA to 150mA by changing The 20V source, by how many volts should you changes the source? To what new value should You set it


1
Expert's answer
2021-12-20T19:58:15-0500

Given:

"I_i= 100mA" "I_f=150mA" , "V_i= 20V"

Resistance "R=\frac{V}{I} =constant"


"\frac{V_i}{I_i} =\frac{V_f}{I_f}"


"\therefore \frac{20}{100} =\frac{V_f}{150}" "V_f=30volts"


The new value should be set at 30 volts